A 0.124 mol sample of nitrogen gas has a volume of 941 milliliters at a pressure of 4.93 atm. The temperature of the gas sample on the Celsius scale is °C.
Answer #1

Given:
P = 4.93 atm
V = 941.0 mL
= (941.0/1000) L
= 0.941 L
n = 0.124 mol

use:
P * V = n*R*T
4.93 atm * 0.941 L = 0.124 mol* 0.08206 atm.L/mol.K * T
T = 456 K
= (456 - 273) oC
= 183 oC
Answer: 183 oC